Let A be 3x3 matrix such that AA=I and |A|=I, find the value of |A-I|


*( means transpose)


 PLZ REPLY SOON...................


Grade:12th pass

1 Answers

Sunil Kumar
9 years ago
HI SKS, BY T, i mean transpose like AT = transpose OF A det (A-I) = det(A-I) . detAT because (det A = det AT = 1) = det(A.AT - AT) because (det AB = det A.detB) = det(I - AT)= - det (AT -I) = -det (A-I)T = -det(A-I) because I = I T AND (X+Y)T = XT + YT = det (A-I) = 0 SUNIL KUMAR IIT K askiitians faculty